OpenStudy (lukecrayonz):

This is confusing me D: Solve: 2x=4+sqrt(7x-12) This is confusing me D: Solve: 2x=4+sqrt(7x-12) @Mathematics

OpenStudy (anonymous):

take the 4 to the left side, then square both sides

OpenStudy (anonymous):

subtract 4, then square both sides

OpenStudy (anonymous):

what lagrange said

OpenStudy (anonymous):

of course when i say "take" i mean subtract

OpenStudy (anonymous):

that was wrong, it should be \[4x^2-16x+16=7x-12\]

OpenStudy (lukecrayonz):

Please fully solve. I did that, I got 4x^2-16x+16=7x-12.

OpenStudy (anonymous):

ok good

OpenStudy (lukecrayonz):

I just need a full explanation, do I factor that now?

OpenStudy (anonymous):

then \[4x^2-23x+28=0\] before you can solve

OpenStudy (anonymous):

now by some miracle this will factor as \[(x-4)(4x-7)=0\] so \[x=4, x=\frac{7}{4}\]

OpenStudy (lukecrayonz):

Well in the back of my textbook, it only has the answer of 4, so why do I have 7/4?

OpenStudy (anonymous):

but you still have to check both answers to see if they work, because you squared

OpenStudy (anonymous):

yes you need to check

OpenStudy (lukecrayonz):

Ahh, alright, sorry, I was absent for this lesson :S

OpenStudy (anonymous):

when you square both sides you can introduce a solution that was not a solution to the original one. for example if i solve \[x-2=4\] i get \[x=6\] but if i square and solve \[(x-2)^2=16\] i get \[x=6,x=-2\] but only one is a solution to the original equation

OpenStudy (lukecrayonz):

I have another question, let me make a new question :)

OpenStudy (anonymous):

ok